A common automotive grade ductile iron casting targets the carbon equivalency to 4.54. If the desired carbon content is 3.55 to 3.70% carbon, what is the range of silicon required to produce this type of ductile iron? How much of this alloy is eutectic?
Answer #1

Answer:

The solution is given below in the fig.

If any query , please comment.
Answer #2

To produce ductile iron with the desired carbon content range of 3.55 to 3.70%, we need to ensure that the carbon equivalency is maintained at 4.54. The carbon equivalency is given by the formula:

CE = %C + (%Si / 4) + (%Mn / 6) - (%P / 20) - (%S / 20)

where %C, %Si, %Mn, %P, and %S are the percentages of carbon, silicon, manganese, phosphorus, and sulfur in the alloy.

Since we want to keep the carbon content between 3.55 and 3.70%, we can assume an average of 3.625% carbon. Substituting the values into the formula, we get:

4.54 = 3.625 + (%Si / 4) + (%Mn / 6) - (%P / 20) - (%S / 20)

Solving for %Si, we get:

%Si = (4.54 - 3.625 - (%Mn / 6) + (%P / 20) + (%S / 20)) * 4

Since the amount of manganese, phosphorus, and sulfur is not given, we can assume that they are present in negligible amounts or within acceptable limits. Therefore, we can simplify the equation to:

%Si = (4.54 - 3.625) * 4 = 3.64

So, the required range of silicon is 3.64%.

To find the amount of eutectic in the alloy, we need to know the composition of the eutectic. In ductile iron, the eutectic is composed of ferrite and graphite. The amount of eutectic is determined by the carbon equivalent of the alloy. For a carbon equivalent of 4.54, the eutectic fraction is approximately 10%. Therefore, 10% of the alloy will be eutectic.


answered by: Hydra Master